### v2

# Anisotropic Ginzburg-Landau model for superconductivity with five-dimensional operators

Abstract
This paper presents the effects of non-minimal Lorentz-violation operators in superconductivity. By constructing a Lorentz-Violating Ginzburg-Landau theory of superconductivity with a five-dimensional operator, we discuss the influence of higher dimensional Lorentz-Violating operators in the London’s depth penetration, in the coherence length and critical magnetic field.
I Introduction
In recent years, there has been a focus on exploring potential expansions of the Standard Model (SM). Within this framework, the discussion has shifted towards the breaking of Lorentz and CPT symmetries, considered as a significant subject Kos1 ; Kos2 ; Kos3 ; Kos4 ; Colladay:1996iz ; Colladay:1998fq . Typically, Lorentz symmetry is disrupted by introducing specific directions in space-time, represented by additive terms that are proportionate to small constant vectors or tensors. The Standard Model Extension (SME) Colladay:1996iz ; Colladay:1998fq stands out as the most established model addressing the consequences of Lorentz and CPT symmetry violation. SME, being an effective field theory, incorporates all conceivable terms violating Lorentz and CPT symmetries in its Lagrangian.
Kostelecky Edwards:2018lsn . proposed an extension of the scalar sector incorporating Lorentz-violating effects. This model presents a comprehensive effective scalar field theory across any spacetime dimension, featuring explicit perturbative spin-independent Lorentz-violating operators of arbitrary mass dimension. The significance of this development lies in the predominance of spin in most fundamental particles of the SM, with the Higgs boson being the sole example of a fundamental spinless particle. Despite the relatively minor role of the scalar sector of QED (sQED) compared to strong interactions in describing meson coupling, it has been suggested Edwards:2019lfb that a Lorentz-violating extension of sQED could effectively address slight CPT deviations in neutral-meson oscillations.
The minimal extension of SME was extensively studied in the last years in several contexts, such as radiative corrections Jackiw:1999yp ; Kostelecky:2001jc ; Ferrari:2021eam , neutrinos oscillations MINOS:2008fnv ; MINOS:2010kat ; MiniBooNE:2011pix , Euler-Heisenberg effective action Ferrari:2021eam ; Furtado:2014cja , gravitational context Kostelecky:2016kkn ; Assuncao:2018jkq , finite temperature field theory Assuncao:2018jkq ; Leite:2013pca ; Leite:2011jg ; Assuncao:2016fko ; Araujo:2023izx among others (for a more complete review, see f.e. Mariz:2022oib ). The non-minimal extension, despite its posses only nonrenormalizable terms, have been receiving attention in the literature Myers:2003fd ; Kostelecky:2009zp ; Kostelecky:2011gq ; Mariz:2010fm ; Rubtsov:2012kb . A strong motivation lies in the fact that some relevant astrophysical processes impose severe restrictions on the coefficients associated with the operators with dimension $d\geq 5$, being such contributions comparable or even dominant when compared with the ones that arise from dimension $d\leq 4$ operators. Within the nonminimal LV framework, an important role was played by the paper Kostelecky:2009zp , in which the simplest cases of such operators were, for the very first time, introduced in the LV scenario. The first studies of their perturbative impact were investigated in Gomes:2009ch ; Mariz:2010fm ; Gazzola:2010vr ; Mariz:2011ed ; BaetaScarpelli:2013rmt . And finally, all possible LV extensions of SME with dimensions up to 6 for fermion-dependent operators were listed in Kostelecky:2018yfa .
The potential applications of models violating Lorentz and CPT symmetries span various fields, extending from quantum field theory to condensed matter physics. The intersection of Lorentz-violating models and condensed matter physics has garnered significant attention, particularly in the context of Weyl semimetals Assuncao:2015lfa ; Gos1 ; Gos2 ; Gos3 , superconductivity Bazeia:2016pra ; Furtado:2020iom ; Granado:2023czl , graphene with anisotropic scaling Katsnelson:2012cz , dark matter and black holes analog models Baym:2020uos ; Pereira:2009vb , among others.
The Ginzburg-Landau (GL) theory of superconductivity serves as a construct within effective quantum field theory, elucidating several crucial facets of superconductivity Ginzburg:1950sr . Despite its initial classification as a phenomenological theory, the GL theory can be construed as a specific case arising from the Bardeen-Cooper-Schrieffer (BCS) theory of superconductivity Bardeen:1957mv . The exploration of superconductivity in diverse realms, such as Weyl semimetals Zhou:2015qka ; Bednik:2015tha ; Wei:2014vsa and carbon-based nanostructures like graphene and fullerene Roy:2013aya ; Cohnitz:2017vsr is frequently undertaken in literature due to their potential applications for technological advancements. However, recent years have seen an intensified discourse on the theoretical aspects of superconductivity, extending into domains such as astrophysics Madsen:1999ci ; Bonanno:2011ch , cosmology Ebert:2007ey ; Gao:2012aw and in high energy physics Herzog:2009xv ; Ghoroku:2019trx . Recently, studies of anisotropic superconductors were carried out by considering Lorentz-violating terms inspired by the SME Furtado:2020iom . The authors have shown in Furtado:2020iom that the coherence length and critical magnetic field are not affected by Lorentz-violating modifications in the scalar sector, but the London’s depth penetration in modified by the Lorentz-violating contributions.
In this study we consider a Lorentz-violating complex scalar sector coupled non-minimally with the gauge field, the usual Maxwell term as the gauge sector and $\lambda|\phi|^{4}$ promoting the spontaneous symmetry breaking. We discuss the influence of a higher dimensional Lorentz-Violating operator in the London’s depth penetration, in the coherence length and critical magnetic field. The Lorentz-violating parameters can be interpreted as defects or layers in the superconductor, giving rise to an anisotropy in the system.
This paper is organized as follows: in the next section we present our model, discussing the discrete symmetries and the non-minimal Lorentz-violating extension of the Ginzburg-Landau theory for superconductivity. In the section III we discuss the superconducting case for the Lorentz-violating contributions to the coherence length, critical magnetic field and London’s depth penetration. Finally, in section IV we highlight our conclusions.
II Model
The model we are considering consists of the complex scalar sector coupled non-minimally with the gauge field, the usual Maxwell term as the gauge sector and $\lambda|\phi|^{4}$ potential promoting the spontaneous symmetry breaking. Hence the Lagrangian describing the system is
$\mathcal{L}$ $=$ $(\nabla_{\mu}\phi)^{*}(\nabla^{\mu}\phi)-m^{2}\phi^{*}\phi-\lambda(\phi^{*}\phi)^{2}-\frac{1}{4}F^{\mu\nu}F_{\mu\nu},$
where
$\nabla_{\mu}\phi=\left(D_{\mu}-i\frac{g}{2}\epsilon_{\mu\nu\alpha\beta}\omega^{\nu}F^{\alpha\beta}\right)\phi$ (1)
is the non-minimal covariant derivative and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. The usual covariant derivative in turn is given by $D_{\mu}=\partial_{\mu}-ieA_{\mu}$. In the high-energy physics scenario, the vector $\omega^{\mu}$ promotes the violation of Lorentz invariance by setting a privileged direction in space-time and breaking the equivalence between particle and observer transformations. Assuming that the vector is constant, it implies the independence of space-time position, which yields translational invariance and assures the conservation of momentum and energy. The interest in this specific model of Lorentz symmetry violation lies in the fact that the covariant derivative in Eq. (1) contains a Pauli-type coupling term, where the complex scalar field directly couples to the electromagnetic field strength. Notice that the product $g\omega^{\mu}$ must have dimension of $(\text{mass})^{-1}$ and it plays the role of a Lorentz violating induced magnetic moment. In $(2+1)$ dimensions, it has been shown that a similar non-minimal coupling term gives rise to fractional spin statistics even in the absence of the Chern-Simons term Nobre:1999mj ; Furtado:2011zz , which is known to exhibit such a characteristic Semenoff:1988jr ; Dunne:1994uy . Fractional statistics play an important role in the interaction between quantum field theory and condensed matter physics Nobre:1999mj ; Wilczek:1982wy .
The kinetic term can be written as
$(\nabla_{\mu}\phi)^{*}(\nabla^{\mu}\phi)$ $=$ $(D_{\mu}\phi)^{*}(D^{\mu}\phi)$
$+$ $i\frac{g}{2}\epsilon_{\mu\nu\alpha\beta}\omega^{\nu}F^{\alpha\beta}\left[\phi^{*}(D^{\mu}\phi)-(D^{\mu}\phi)^{*}\phi\right]+\cdots,$
where the dots stands for higher-orders Lorentz breaking terms.
In the static condition the Lagrangian takes the form:
$\mathcal{L}$ $=$ $\eta^{ij}(D_{i}\phi)^{*}(D_{j}\phi)-\mu^{2}\phi^{*}\phi-\lambda(\phi^{*}\phi)^{2}-\frac{1}{4}F_{ij}F^{ij}$ (3)
-- $i\,\frac{g}{2}W\epsilon_{ijk}F^{jk}\left[\phi^{*}D^{i}\phi-(D^{i}\phi)^{*}\phi\right].$
Now, $-\mathcal{L}$ is the anisotropic Ginzburg-Landau free energy with a five-dimensional operator. Note that the mass parameter $m^{2}$ was written as $\mu^{2}$, which is now considered a temperature-dependent system parameter near the critical temperature $T=T_{c}$. Specifically, $\mu^{2}=a\,(T-T_{c})$. We have further identified the temporal component of the vector $\omega^{\mu}$ as being $W$, i.e., $W=\omega^{0}$. In this context, $\phi$ is the macroscopic many-particle wave function. The physical interpretation of $\phi$ as a many-particle wave function is justified by the Bardeen-Cooper-Schrieffer (BCS) theory, according to which, under certain conditions, there is an attractive force between electrons, and field quanta are electron pairs, which are, of course, bosons. At low temperatures, the field quanta fall into the same quantum state (Bose-Einstein Condensation Furtado:2020olp ; Casana:2011bv ), and because of this, a many-particle wave function $\phi$ may be used to describe the macroscopic system. In this sense, the quantity $e$ in the usual covariant derivative expression should be understood as being twice the value of the electric charge.
The equation of motion for the gauge field is given by
$\mathbf{\nabla}^{2}A^{m}+\partial_{n}\partial^{m}A^{n}=-ej^{m},$ (4)
where
$j^{m}$ $=$ $-i\phi^{*}\overleftrightarrow{\partial^{m}}\phi-2eA^{m}\phi^{*}\phi+gW(\mathbf{x})\epsilon^{mjk}F_{jk}\phi^{*}\phi$ (5)
$+$ $i\,\frac{g}{e}\partial_{n}W(\mathbf{x})\epsilon^{imn}(\phi^{*}\overleftrightarrow{\partial_{i}}\phi-2ieA_{i}\phi^{*}\phi)$
$+$ $i\,\frac{g}{e}W(\mathbf{x})\epsilon^{imn}\Bigl{[}2\partial_{n}\phi^{*}\partial_{i}\phi+\phi^{*}\overleftrightarrow{\partial_{ni}}\phi$
-- $2ie\partial_{n}A_{i}\phi^{*}\phi-2ieA_{i}(\phi^{*}\partial_{n}\phi+\partial_{n}\phi^{*}\phi)\Bigr{]}$
is the conserved current density due to the invariance of the theory under global $U(1)$ symmetry. In equation (5),
$\phi^{*}\overleftrightarrow{\partial_{i_{1},\cdots,i_{n}}}\phi=\phi^{*}(\partial_{i_{1}}\cdots\partial_{i_{n}})\phi-(\partial_{i_{1}}\cdots\partial_{i_{n}})\phi^{*}\phi,$ (6)
and we have taken into account the possibility that the Lorentz violation parameter is a function that depends on position, i.e. $W=W(\mathbf{x})$.
Considering that $\phi$ varies only very slightly over the sample, so that $\partial_{i}\phi\approx 0$, we can simplify the current density $j^{m}=j^{m}(\mathbf{x})$ to
$j^{m}(\mathbf{x})$ $=$ $|\phi|^{2}\Bigl{[}-2eA^{m}+gW(\mathbf{x})\epsilon^{mjk}F_{jk}$ (7)
$+$ $2g\epsilon^{imn}\partial_{n}W(\mathbf{x})A_{i}+2g\epsilon^{imn}W(\mathbf{x})\partial_{n}A_{i}\Bigr{]}$
Now, the ground state where Bose-Einstein Condensation is conceived can be obtained by minimizing the potential $V(\phi^{*},\phi)$ for a temperature $T<T_{c}$, i.e., $\mu^{2}<0$. Since
$V(\phi^{*},\phi)=\mu^{2}|\phi|^{2}+\lambda|\phi|^{4},$ (8)
the free energy minimum occurs at
$|\phi|^{2}=-\frac{\mu^{2}}{2\lambda}>0,$ (9)
where, by definition, $\lambda>0$. In this configuration, the current density in Eq. (7) takes the form of
$\mathbf{j}(\mathbf{x})=-\Gamma^{2}\mathbf{A}-\frac{g}{e}\Gamma^{2}\Biggl{[}2W(\mathbf{x})(\mathbf{\nabla}\times\mathbf{A})+\mathbf{\nabla}W(\mathbf{x})\times\mathbf{A}\Biggr{]}$ (10)
where
$\Gamma^{2}\equiv-\frac{e\mu^{2}}{\lambda},$ (11)
with $\Gamma$ being a positive definite constant. From now on, we will only work with the case when $T<T_{c}$. Therefore, no phase transitions will be considered. Eq. (10) is the so called London equation, modified by the presence of the Lorentz-violating function $W(\mathbf{x})$.
Now, Ohm’s law defines resistance $R$ through the equation $\mathbf{E}=R\,\mathbf{j}$. In the static condition of our model, the electric field is simply $\mathbf{E}=-\partial\mathbf{A}/\partial t=\mathbf{0}$. As a result, we have two possible solutions: $R=0$ (superconducting case) or $\mathbf{j}=\mathbf{0}$ (insulating case).
A non-trivial solution for the insulating case occurs when $W(\mathbf{x})\neq 0$. To ensure that the current is zero within the insulating material, the vector potential and the Lorentz violation parameter must satisfy
$\mathbf{A}+\frac{g}{e}\,\Biggl{[}2W(\mathbf{x})(\mathbf{\nabla}\times\mathbf{A})+\mathbf{\nabla}W(\mathbf{x})\times\mathbf{A}\Biggr{]}=\mathbf{0}.$ (12)
From Faraday’s law we know that the magnetic field must also be static and since Ampère-Maxwell equation leads to ×$\mathbf{\nabla}\times\mathbf{B}=\mathbf{0}$, taking the curl from Eq. (12), we find
$2\,\mathbf{B}\times\mathbf{\nabla}W(\mathbf{x})-\frac{e}{g}\,\mathbf{B}$ $=$ $\bigl{(}\mathbf{A}\cdot\mathbf{\nabla}\bigr{)}\,\mathbf{\nabla}W(\mathbf{x})-\bigl{(}\mathbf{\nabla}W(\mathbf{x})\times\mathbf{\nabla}\bigr{)}\,\mathbf{A}$ (13)
$+$ $\mathbf{\nabla}W(\mathbf{x})\,(\mathbf{\nabla}\cdot\mathbf{A})-\mathbf{A}\,\mathbf{\nabla}^{2}W(\mathbf{x}).$
As we can notice from Eq. (13), the magnetic field will be zero if the violation parameter remains constant over the material.
III Superconducting case
In this section we will study the superconducting case ($R=0$) for both constant and position-dependent Lorentz-violating parameter. In the context of GL theory of superconductivity, three important properties can be obtained, namely, the coherence length, the critical magnetic field, and London’s penetration depth. Let’s investigate each of these properties separately in the following.
III.1 The coherence length
In order to obtain the coherence length of the superconductor we have to solve the equation of motion for the complex scalar field, which is given by
$\Lambda\phi-2\lambda\,|\phi|^{2}\phi$ $=$ $ig\,\Bigl{\{}2W(\mathbf{x})\,\bigl{[}(\mathbf{\nabla}\times\mathbf{A})\cdot\mathbf{D}\bigr{]}$ (14)
-- [(×$\bigl{[}(\mathbf{\nabla}\times\mathbf{A})\cdot\mathbf{\nabla}W(\mathbf{x})\bigr{]}\Bigr{\}}\phi$
where we have defined $\Lambda$ as
$\Lambda\equiv\mathbf{\nabla}^{2}-\mu^{2}+2ie\,\mathbf{A}\cdot\mathbf{\nabla}+ie\,\mathbf{\nabla}\cdot\mathbf{A}-e^{2}\,\mathbf{A}^{2},$ (15)
and used the fact that $\mathbf{D}=-\mathbf{\nabla}-ie\mathbf{A}$. However, the coherence length is a property that characterizes superconducting materials and is independent of the existence of an external magnetic field. For that reason we can assume $\mathbf{A}=\mathbf{0}$, since $\mathbf{A}$ is the vector potential that arises within the material in response to the external magnetic field. Regarding boundary conditions, we usually desire the field $\phi$ to be zero at the surface of the superconductor and acquire a nearly constant maximum value of $\phi_{max}$ deep inside the material. Note from Eq. (14) that $|\phi_{max}|^{2}=-\mu^{2}/2\lambda$. Based on what was discussed above, it is straightforward to show that the superconducting electrons density must be given by
$|\phi|^{2}=|\phi_{max}|^{2}\tanh^{2}\left(\frac{x}{\sqrt{2}\,\xi}\right)$ (16)
for a semi-infinite superconductor occupying the space $x>0$. The parameter $\xi$, given by
$\xi=\sqrt{\frac{1}{|\,\mu^{2}\,|}},$ (17)
defines the coherence length. Note that this parameter is insensitive to Lorentz-violating effects.
III.2 The critical magnetic field
The critical magnetic field is defined as the minimum value of the external magnetic field $H$ above which the state of superconductivity is lost. For a type-I superconductor, the critical magnetic field $H_{c}$ is directly calculated from the Ginzburg-Landau free energy under the condition that it is the maximum value of the external magnetic field for which the free energy of the superconducting state is less than that of the normal state. In this sense, in the presence of an external magnetic field and at the limit where superconductivity is maximum, we can write
$\mu^{2}\,|\phi_{max}|^{2}+\lambda\,|\phi_{max}|^{4}+\mu_{0}\,\frac{H_{c}^{2}}{2}=0$ (18)
so that
$H_{c}=\frac{|\mu^{2}|}{\sqrt{2\lambda\mu_{0}}}.$ (19)
Note from Eq. (19) that the critical magnetic field for a type-I superconductor is not affected by the presence of Lorentz violation in our model.
In type-II superconductors, the value of $H_{c}$ is not sufficient to break the superconducting state, since the material can exist in a vortex state. To determine the critical magnetic field $H_{c2}$ above which the normal state of a type-II superconducting material is reached, we must analyze the equation of motion for the complex scalar field in a specific geometry. Let us consider a semi-infinite type-II superconductor occupying the space $x>0$ under an external magnetic field $\mathbf{H}=H\hat{\mathbf{z}}$ with $H$ constant. Since magnetization is small, especially in strongly type-II superconductors, we can assume that inside the superconductor $B\approx\mu_{0}H$. Therefore, we can suppose a vector potential of the form
$\mathbf{A}=xB\hat{\mathbf{y}},$ (20)
which is consistent with the Landau gauge ($\mathbf{\nabla}\cdot\mathbf{A}=0$). Also, for the sake of simplicity, we will assume that the Lorentz violation parameter is a constant $W$. Thus, since it is sufficient to consider the Lorentz-violating linearized Ginzburg-Landau equation at the threshold of superconductivity, Eq. (14) takes the form of
$\Bigl{(}\mathbf{\nabla}^{2}-\mu^{2}+2iexB\partial_{y}-2^{2}x^{2}B^{2}+2igWB\partial_{z}\Bigr{)}\,\phi=0,$ (21)
where $\partial_{a}=\partial/\partial a$. Due to the boundary conditions on the surface and interior of the superconductor, which were mentioned earlier, we are led to consider solutions of the form
$\phi(\mathbf{x})=\varphi(x)\,e^{i(k_{y}y+k_{z}z)}.$ (22)
Substituting Eq. (22) into Eq. (21), we obtain
$\Biggl{[}-\frac{d^{2}}{dX^{2}}+(eBX)^{2}-(|\mu^{2}|-k_{z}^{2}-2gWBk_{z})\Biggr{]}\,\varphi=0,$ (23)
where we have made the displacement $x=X-(k_{y}/eB)$. Eq. (23) can then be identified as the Schrödinger equation for a harmonic oscillator of mass $m=1/2$, frequency $\omega=2eB$, and energy eigenvalue $E=|\mu^{2}|-k_{z}^{2}-2gWBk_{z}$. As we know, the solution to this equation is given in terms of Hermite polynomials, which are stationary and normalized solutions. However, there exists a limiting value of $B$ for which this type of solution can exist. This value, which is related to the maximum value that the potential $V(X)=(1/4)\omega^{2}X^{2}$ can have for the Schrödinger equation to have a solution, occurs when the energy eigenvalue is at its minimum, i.e., $E_{0}=\omega/2$. Thus, we have
$B_{\text{max}}=\mu_{0}H_{\text{max}}=\frac{|\mu^{2}|-k_{z}^{2}}{e+2gWk_{z}}.$ (24)
The critical external magnetic field $H_{c2}$ is defined for the value of $k_{z}$ that makes the above expression a maximum, that is:
$H_{c2}=\frac{2e-2\sqrt{e^{2}-4|\mu^{2}|g^{2}W^{2}}}{4\mu_{0}g^{2}W^{2}}.$ (25)
Unlike what we find in equation (19), our result for the critical magnetic field $H_{c2}$, capable of destroying the superconductivity state of a type II superconductor, explicitly feels the effects of Lorentz violation characterized by the constant parameter $W$. Since the Lorentz-violating effects are expected to be very small, we can still approximate Eq. (25) as
$H_{c2}\approx\frac{|\mu^{2}|}{\mu_{0}e}+\frac{g^{2}W^{2}|\mu^{2}|^{2}}{\mu_{0}e^{3}}.$ (26)
We can then identify the correction to the usual critical magnetic field as $\delta H_{c2}=g^{2}W^{2}|\mu^{2}|^{2}/\mu_{0}e^{3}$. Note that this correction is intended to increase the value of $H_{c2}$, as we can see from Fig. (1). Also note that the linear behavior found in the usual case (black line) is in accordance with what is expected from the Ginzburg-Landau theory for superconductors in the vicinity of $T=T_{c}$.
Figure 1: Graph of the critical magnetic field $H_{c2}$ as a function of the temperature $T$ for different values of the product $gW$. We have considered $a=6.3\times 10^{-19}$ and $T_{c}=18.05\,K$ (Critical temperature of the type II superconductor $Nb_{3}Sn$).
III.3 London’s penetration depth
In order to investigate the influence of Lorentz violation on the London’s penetration depth, we must turn our attention to the current density found in Eq. (10). Taking the curl of both sides of it and making use of Ampère-Maxwell’s law, we can write
$\mathbf{\nabla}^{2}\mathbf{B}$ $=$ $\Gamma^{2}\mathbf{B}+\frac{g}{e}\Gamma^{2}\Biggl{\{}2W(\mathbf{x})\bigl{[}\mathbf{\nabla}\times\mathbf{B}\bigr{]}-2\mathbf{B}\times\mathbf{\nabla}W(\mathbf{x})$ (27)
$+$ $\bigl{[}\mathbf{A}\cdot\mathbf{\nabla}\bigr{]}\mathbf{\nabla}W(\mathbf{x})-\bigl{[}\mathbf{\nabla}W(\mathbf{x})\cdot\mathbf{\nabla}\bigr{]}\mathbf{A}$
$+$ $\mathbf{\nabla}W(\mathbf{x})\bigl{[}\mathbf{\nabla}\cdot\mathbf{A}\bigr{]}-\mathbf{A}\mathbf{\nabla}^{2}W(\mathbf{x})\Biggr{\}}.$
From the above equation, in the absence of any violation, we recover the result that defines the usual London penetration depth $\lambda_{L}$, namely
$\mathbf{\nabla}^{2}\mathbf{B}=\frac{1}{\lambda_{L}^{2}}\,\mathbf{B}$ (28)
where $\lambda_{L}=\Gamma^{-1}=\sqrt{\lambda/e|\mu^{2}|}$. A much simpler form of Eq. (27) is obtained when the violation parameter is a constant $W$:
$\mathbf{\nabla}^{2}\mathbf{B}$ $=$ $\frac{1}{\lambda_{L}^{2}}\,\left[\mathbf{B}+\left(\frac{2gW}{e}\right)\,\mathbf{\nabla}\times\mathbf{B}\right]$ (29)
IV Final Remarks
In this paper we study the effects of non-minimal Lorentz-violation operators in superconductivity. By constructing a Lorentz-Violating Ginzburg-Landau theory of superconductivity with five-dimensional operators, we discuss the influence of a higher dimensional Lorentz-Violating operator in the London’s depth penetration, in the coherence length and critical magnetic field that breaks the superconducting state.
The higher-dimensional Lorentz-violating operator was included by considering a non-minimal coupling between the complex scalar field, which stands as the field for the Cooper pairs, and the gauge field. We shown that the coherence length is not affected by the Lorentz-violating contributions.
Also, for type I superconductors, we have demonstrated that the critical magnetic field is insensitive to the Lorentz-violating contributions. However, the critical magnetic field $H_{c2}$ for type II superconductors is affected by the Lorentz-violating contributions, so that the increasing in the LV term $gW$ leads to an increase in the value of $H_{c2}$.
The London’s depth penetration is drastically modified by the presence of the Lorentz-violating contributions. Here we have considered two cases, namely, the constant $W$ parameter case and a position-dependent $W(\bf{x})$ case. For both cases we have depicted the behaviour of the magnetic field inside the superconductor in order to investigate the influence of the Lorentz-violating term in the Meissner effect.
Finally, it was evident that under certain conditions controlled by the Lorentz-violating contribution, there exists a first-order phase transition between the superconducting state and an insulating state. It is important to highlight here that the Lorentz-violating parameter $W$ plays the role of local impurities that could break the homogeneity and isotropy of the system as a whole. Also, it is worth to mention that the addition of new terms violating Lorentz symmetry, whether in the pure scalar sector Furtado:2020iom , or through different couplings between the scalar sector and the gauge sector or even direct modifications in the pure gauge sector, as proposed by other types of electrodynamics, should lead to results completely different from those presented here. Such additional terms directly impact the equations for the scalar and gauge fields, from which the quantities calculated here are derived. Specifically, we expect modifications both in the critical magnetic field $H_{c2}$ for a type II superconductor and in the behavior of the magnetic field within the superconductor, i.e., related to the London penetration depth. In fact, this will be a subject of investigation in future works.
Acknowledgements.
      JF would like to thank the Fundação Cearense de Apoio ao Desenvolvimento Científico e Tecnológico (FUNCAP) under the grant PRONEM PNE0112-00085.01.00/16 for financial support and Gazi University for the kind hospitality.